JoyLoveReynolds:

Find the polynomial. 2(3y − 1)(y − 3) is the factored form of

JoyLoveReynolds:

ThisGirlPretty:

@theDeviliscoming

theDeviliscoming:

hold up lemme get scratch paper

theDeviliscoming:

2(3y − 1)(y − 3) = 2(3y)(y) - 2(y) - 2(3y) (3) + 2(1)(3) = 6y^2 -2y - 18y + 6 = 6y^2 - 20y + 6